When PCL5 is heated it gasifies and dissociates into PCL3 and CL2. The density of the gas mixture at 200°C is 70.2 find the degree of dissociation of PCL5 at 200°C.

Grade:12

1 Answers

Arun
6 years ago
initial vapour densuty/vapour density at equilibrium=mole at equilibrium/mole initialPCL5 -----PCL3 +CL2let the initial concentration be Cconcentration of PCL3=C-Cx,x is the degree of dissociationPCL3=CxCL2=Cxtotal mole at equilibrium=C +Cxinitial mole=CthereforeC +Cx/C =104.25/701+x=1.485x=.485